Grade 12th passZoology

A train is traveling at a speed of 90 km h-1. Brakes are applied so as to produce a uniform acceleration of? 0.5 m s-2. Find how far the train will go before it is brought to rest.

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Initial speed of the train, u= 90 km/h = 25 m/s (1km/hr = 5/18 m/s)
Final speed of the train, v = 0 (finally the train comes to rest and its velocity becomes 0)
Acceleration = - 0.5 ms-2
According to third equation of motion:
v2= u2 + 2 as
(0)2= (25)2 + 2 ( - 0.5) s
Where, s is the distance covered by the train
S = 25​2
2 x 0.5
= 625m

Thanks